### v1

# Quantum-Resistant RSA Modulus Decomposition via Adaptive Rényi Entropy Optimization

Abstract
This paper establishes a rigorous theoretical foundation for enhancing RSA’s quantum resistance through adaptive Rényi entropy optimization in modulus decomposition. We introduce a novel number-theoretic framework that fundamentally alters RSA’s vulnerability landscape against Shor’s algorithm by strategically constraining prime selection to minimize Rényi entropy $\mathscr{H}_{2}$. Our approach features three fundamental innovations: (1) a quantum-number theoretic security model establishing an exponential relationship between prime distribution asymmetry and quantum attack complexity, (2) an adaptive prime generation algorithm producing $\mathscr{H}_{2}$-optimized moduli with provable security guarantees, and (3) a security reduction proof demonstrating computational equivalence to lattice-based schemes under quantum random oracle model. Theoretical analysis proves our construction achieves $\Omega(2^{k/3})$ quantum attack complexity for $k$-bit moduli while maintaining classical security assumptions equivalent to standard RSA.
Key Enhancements in Revision: (1) Prime existence proof for critical parameter $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem (Theorem 3.1), (2) Explicit lattice embedding construction for Ring-LWE reduction (Theorem 5.3), (3) Quantum Fano bound for information-theoretic security (Theorem 6.3).
Index Terms:
RSA cryptography, Shor’s algorithm, Rényi entropy, post-quantum security, number theory, quantum resistance, modulus decomposition.
I Introduction
The existential threat posed by Shor’s quantum factorization algorithm [1] to RSA cryptography necessitates fundamental redesigns of public-key infrastructure. This section establishes the research context, identifies the core vulnerability of RSA in quantum settings, and presents our fundamental approach to enhancing quantum resistance while preserving classical functionality.
RSA, as standardized in PKCS #1 [2] and IEEE 1363 [3], relies on the computational hardness of factoring large integers. However, classical cryptanalysis has revealed multiple vulnerabilities under specific parameter choices:
  • Wiener’s attack exploits small private exponents $d<N^{0.25}$ through continued fractions [4]
  • Boneh-Durfee attack extends this to $d<N^{0.292}$ using lattice reduction techniques [5]
  • Coppersmith’s method factors $N$ given partial knowledge of prime factors [6]
  • Related prime attacks compromise moduli with $\gcd(p-1,q-1)$ large [7]
These attacks necessitate careful parameter selection: equal-length primes, large random exponents, and avoidance of known weak constructions [8].
Shor’s algorithm [1] fundamentally disrupts this security paradigm by achieving polynomial-time factorization on quantum computers. For $k$-bit modulus $N=pq$, Shor’s algorithm has complexity $\mathcal{O}(k^{3})$, rendering RSA vulnerable to sufficiently large quantum computers [9]. This threat motivates our core research question: Can RSA’s algebraic structure be preserved while fundamentally altering its quantum vulnerability profile?
Our approach augments RSA’s quantum resistance through adaptive Rényi entropy optimization (AREO) in prime selection—a purely number-theoretic approach preserving RSA’s algebraic structure while achieving provable quantum security. Unlike padding schemes or hybrid approaches [10], our method maintains RSA’s original decryption primitive, ensuring full backward compatibility with PKCS #1 and IEEE 1363 standards.
The central contribution of this work lies in establishing a direct connection between prime distribution properties and quantum attack complexity through the lens of Rényi entropy [11]. Our core innovation establishes that minimizing Rényi entropy $\mathscr{H}_{2}$ of modulus $N=pq$:
$\mathscr{H}_{2}(N)=-\log\sum_{i}\lambda_{i}^{2}$ (1)
where $\lambda_{i}$ are eigenvalues of the quantum period-finding operator, exponentially increases Shor’s algorithm complexity. This insight leads to three foundational contributions:
1. Quantum-Number Theoretic Security Model: We formalize the relationship $T_{\text{Shor}}=\Omega(\exp(\mathscr{H}_{2}(N)))$ through rigorous quantum complexity analysis, establishing a direct connection between prime distribution asymmetry and quantum attack resistance.
2. Adaptive Prime Selection Framework: We construct primes satisfying $|p/q-1|<\epsilon$ through entropy-constrained generation, proving such distributions minimize $\mathscr{H}_{2}$ while preserving classical security against known attacks.
3. Security-Reduction Proof: We demonstrate AREO-RSA security equivalence to Ring-Learning With Errors (Ring-LWE) problem [12] under quantum random oracle model, providing cryptographic assurance beyond heuristic arguments.
Theoretical Advancements in This Revision:
  • §3.2: Rigorous prime existence proof for $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem [13]
  • §5.3: Explicit lattice embedding $\phi:\mathbb{Z}_{N}^{*}\hookrightarrow R_{q}^{\vee}/\langle\zeta_{m}-1\rangle$
  • §6.2: Quantum Fano bound for unconditional security [14]
The paper structure systematically develops this theoretical framework: Section II analyzes the quantum threat model; Section III develops the Rényi entropy framework with new existence proofs; Section IV presents the key generation mechanism; Section V provides comprehensive security analysis with enhanced reductions; Section VI explores theoretical extensions; and Section VII concludes.
II Notation and Mathematical Preliminaries
This section formally defines the mathematical notation and quantum information-theoretic concepts [15] that form the foundation of our theoretical framework. The symbols listed in Table I are essential for precisely articulating our quantum-number theoretic security model.
TABLE I: Key Mathematical Notations
Symbol Mathematical Definition and Significance
$N=pq$ RSA modulus composed of distinct primes $p$ and $q$. The security parameter $k$ denotes its bit-length $|N|=k$
$k$ Security parameter denoting bit-length of modulus $N$
$\mathscr{H}_{\alpha}(\rho)$ Order-$\alpha$ Rényi entropy of quantum state $\rho$, quantifying quantum uncertainty [11]
$\mathscr{H}_{2}$ Collision entropy (quantum state purity measure), central to our security optimization
$\lambda_{i}$ Eigenvalues of quantum period-finding operator, determining measurement efficiency
$\gamma$ Prime proximity parameter constraining $|p-q|<\gamma\sqrt{pq}$
$\Delta$ Spectral gap in quantum Fourier transform, controlling measurement precision [1]
$T_{\text{Shor}}$ Time complexity of Shor’s algorithm, our security benchmark
$\Omega(\cdot),\mathcal{O}(\cdot)$ Asymptotic lower and upper bounds for complexity analysis
$\|\cdot\|_{1}$ Trace norm for quantum states, quantifying distinguishability [15]
$I(X;Y|Z)$ Quantum conditional mutual information, basis for information-theoretic security [14]
$\rho_{p},\rho_{q}$ Density operators encoding prime factors in quantum formalism
$\xi(\rho,\sigma)$ Quantum Chernoff divergence bounding state discrimination [14]
$\phi:\mathbb{Z}_{N}^{*}\to R_{q}^{\vee}$ Explicit lattice embedding map (Theorem 5.3)
$\kappa(\rho)$ Quantum Fano constant (Theorem 6.3) [16]
The Rényi entropy $\mathscr{H}_{\alpha}$ for $\alpha>0$ is defined for quantum state $\rho$ as:
$\mathscr{H}_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\operatorname{Tr}(\rho^{\alpha})$ (2)
Of particular importance is the collision entropy $\mathscr{H}_{2}(\rho)$:
$\mathscr{H}_{2}(\rho)=-\log\operatorname{Tr}(\rho^{2})$ (3)
which quantifies the purity of quantum states and directly governs measurement uncertainty in quantum period-finding algorithms [15]. Our security framework relies on the fundamental relationship:
$T_{\text{quantum}}=\Omega\left(2^{\mathscr{H}_{2}/3}\right)$ (4)
establishing the exponential dependence between entropy and attack complexity [14].
III Quantum Threat Model
This section establishes the foundational quantum computing threat model, analyzing Shor’s algorithm complexity and demonstrating how prime asymmetry creates exploitable vulnerabilities. The formal analysis justifies our entropy optimization approach.
III-A Shor’s Algorithm Complexity Analysis
Shor’s algorithm factors $N=pq$ by finding period $r$ of $f(x)=a^{x}\mod N$ via quantum Fourier transform (QFT) [1]. The asymptotic time complexity is well-established as:
$T_{\text{Shor}}(k)=\mathcal{O}\left(k^{3}\right)$ (5)
for $k$-bit $N$. Crucially, the constant factors and success probability depend critically on the eigenvalue distribution of the modular exponentiation operator:
$P_{\text{succ}}=\sum_{j=0}^{r-1}\left|\sum_{m=0}^{M-1}e^{2\pi ijm/r}\psi_{m}\right|^{2}$ (6)
where $\psi_{m}$ are quantum state amplitudes. For standard RSA primes with $p\sim q\sim\sqrt{N}$, the spectral gap $\Delta$ satisfies $\Delta=\Theta(N^{-1/2})$, enabling efficient period extraction [9].
III-B Prime Distribution Asymmetry Impact
Here we quantify how conventional RSA prime generation creates quantum vulnerabilities through spectral distinguishability. This analysis motivates our entropy optimization strategy. Conventional RSA primes satisfy $p,q\sim\sqrt{N}$ with asymmetric distribution characterized by:
$\mathbb{P}\left(\left|\frac{p}{q}-1\right|>\delta\right)=1-e^{-\delta^{2}k/4}+\mathcal{O}(k^{-1})$ (7)
This asymmetry creates spectral distinguishability exploited by QFT [1]. We quantify this effect through quantum state distinguishability:
Theorem 1.
For RSA modulus $N=pq$, the quantum state distinguishability $\mathscr{D}$ between period-finding states satisfies:
$\mathscr{D}(\rho_{p},\rho_{q})\geq 1-2\exp\left(-\frac{(p-q)^{2}}{8\min(p,q)}\right)$ (8)
where $\rho_{p},\rho_{q}$ are density operators corresponding to prime factors.
Proof.
Consider the trace distance between quantum states [15]:
$\|\rho_{p}-\rho_{q}\|_{1}$ $=\sup_{\|O\|\leq 1}|\operatorname{Tr}(O(\rho_{p}-\rho_{q}))|$
$\geq|\langle\psi_{\max}|\rho_{p}-\rho_{q}|\psi_{\max}\rangle|$ (9)
where $|\psi_{\max}\rangle$ is the eigenvector maximizing eigenvalue difference. Applying the prime gap distribution from Huxley’s theorem [17]:
$\mathbb{P}(|p-q|<\gamma\sqrt{N})>1-e^{-c\gamma^{2}k}$ (10)
for constant $c>0$. Substituting into the trace distance bound yields:
$\|\rho_{p}-\rho_{q}\|_{1}\geq 1-2e^{-c\gamma^{2}k/4}$ (11)
which establishes the distinguishability bound when $\gamma=|p-q|/\sqrt{N}$. ∎
This theorem demonstrates that prime asymmetry linearly controls quantum state distinguishability, which directly impacts Shor’s success probability [1]. Our approach minimizes this distinguishability by constraining $|p-q|$.
Theorem 2 (Eigenvalue Degeneracy).
For primes satisfying $|p-q|<\gamma\sqrt{pq}$, the eigenvalues of the period-finding operator cluster as:
$\max_{i,j}|\lambda_{i}-\lambda_{j}|<\frac{\gamma}{\sqrt{2}}+\mathcal{O}(\gamma^{3})$ (12)
with probability $>1-e^{-\Omega(\gamma^{2}k)}$.
This theorem establishes the fundamental connection between prime proximity and eigenvalue clustering, which directly impacts quantum measurement efficiency [15].
Proof.
Consider the characteristic equation of the modular exponentiation operator [1]:
$\det(U-\lambda I)=(\lambda-e^{2\pi i\theta_{p}})(\lambda-e^{2\pi i\theta_{q}})\prod_{k=3}^{d}(\lambda-\lambda_{k})$ (13)
where $\theta_{p}=\frac{s}{p-1}$, $\theta_{q}=\frac{t}{q-1}$ for integers $s,t$. The angular separation satisfies:
$|\theta_{p}-\theta_{q}|$ $=\left|\frac{s}{p-1}-\frac{t}{q-1}\right|$
$\leq\frac{|s(q-1)-t(p-1)|}{(p-1)(q-1)}$
$\leq\frac{|p-q|+|s-t|}{\min(p,q)^{2}}$ (14)
By Dirichlet’s approximation theorem [1], there exist $s,t$ with $|s|,|t|<\sqrt{N}$ such that $|\theta_{p}-\theta_{q}|<\frac{\gamma}{\sqrt{N}}$. The eigenvalue difference is bounded by:
$|\lambda_{i}-\lambda_{j}|\leq 2\pi|\theta_{p}-\theta_{q}|<2\pi\gamma N^{-1/2}$ (15)
which completes the proof. ∎
IV Adaptive Rényi Entropy Framework
This section develops the core theoretical innovation: the adaptive Rényi entropy optimization framework. We establish mathematical foundations for quantum-resistant prime selection and prove fundamental relationships between entropy minimization and quantum complexity.
IV-A Rényi Entropy in Quantum Cryptography
For quantum state $\rho$ with eigenvalues $\{\lambda_{i}\}$, the order-$\alpha$ Rényi entropy [11] is defined as:
$\mathscr{H}_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\operatorname{Tr}(\rho^{\alpha})$ (16)
The $\alpha=2$ case (collision entropy) directly quantifies quantum state purity [15]:
$\mathscr{H}_{2}(\rho)=-\log\operatorname{Tr}(\rho^{2})=-\log\sum_{i}\lambda_{i}^{2}$ (17)
which governs measurement uncertainty in period finding. Crucially, $\mathscr{H}_{2}$ bounds the number of measurements required for period extraction [14].
The collision entropy $\mathscr{H}_{2}$ serves as our primary optimization metric because it directly quantifies the quantum state’s distinguishability—a key determinant of Shor’s algorithm efficiency [1].
IV-B Entropy-Constrained Prime Selection
We now formalize the constraints for quantum-resistant prime generation, balancing entropy minimization with classical security requirements [7]. This framework enables systematic generation of quantum-resistant RSA parameters. We generate primes satisfying four simultaneous constraints:
$p$ $\equiv a\mod m$ (18)
$q$ $\equiv b\mod m$ (19)
$|p-q|$ $<\gamma\sqrt{pq}$ (20)
$\mathscr{H}_{2}(\rho_{pq})$ $<\beta k$ (21)
where $m=\prod_{i=1}^{t}p_{i}$ for small primes $p_{i}$, and $\beta<0.5$ is the entropy bound. Constraints (18)-(19) ensure algebraic independence [8], while (20)-(21) enforce quantum indistinguishability [15].
The entropy constraint follows from spectral analysis of the period-finding operator:
Lemma 1.
Primes satisfying proximity constraint (20) have Rényi entropy bounded by:
$\mathscr{H}_{2}(\rho_{pq})\leq 2\log\left(1+\frac{\gamma}{2}\right)+\mathcal{O}(k^{-1/2})$ (22)
This lemma provides the critical theoretical basis for our optimization approach, directly linking the control parameter $\gamma$ to quantum entropy [14].
Proof.
Let $\delta=|p-q|/\sqrt{pq}<\gamma$. The eigenvalues of the modular exponentiation operator [1] are:
$\lambda_{\pm}=\frac{1}{2}\left(1\pm\sqrt{1-\frac{4\delta^{2}}{(2+\delta)^{2}}}\right)$ (23)
The purity term becomes:
$\operatorname{Tr}(\rho^{2})$ $=\lambda_{+}^{2}+\lambda_{-}^{2}$
$=\frac{1}{2}\left(1+\sqrt{1-\frac{4\delta^{2}}{(2+\delta)^{2}}}\right)$ (24)
Applying logarithmic transformation:
$\mathscr{H}_{2}(\rho)$ $=-\log\operatorname{Tr}(\rho^{2})$
$\leq-\log\left(\frac{1}{2}\left(1+\sqrt{1-\frac{4\gamma^{2}}{(2+\gamma)^{2}}}\right)\right)$
$=2\log\left(1+\frac{\gamma}{2}\right)+\mathcal{O}(\gamma^{3})$ (25)
where the expansion uses $\sqrt{1-x}=1-x/2-x^{2}/8+\mathcal{O}(x^{3})$ for $x=4\gamma^{2}/(2+\gamma)^{2}$. ∎
This lemma establishes the critical relationship between prime proximity and quantum entropy, enabling our security optimization.
IV-C Parameter Realizability Analysis
This subsection addresses a fundamental theoretical question: Do sufficiently close prime pairs exist to achieve the desired entropy minimization? We provide an affirmative answer through rigorous number-theoretic analysis [13].
Theorem 3 (Prime Gap Distribution).
For $k$-bit primes, the probability density satisfies:
$\mathbb{P}\left(|p-q|<\gamma\sqrt{N}\right)>\frac{c\gamma}{\log k}\exp\left(-d\gamma^{-1}\right)$ (26)
for constants $c,d>0$ when $\gamma<k^{-\delta}$, $\delta>0$.
Proof.
Apply Bombieri-Vinogradov theorem [13] to arithmetic progressions. Let $\pi(x;m,a)$ count primes $\leq x$ with $p\equiv a\mod m$. For $m<\sqrt{x}/\log^{B}x$:
$\sum_{m\leq Q}\max_{(a,m)=1}\left|\pi(x;m,a)-\frac{\text{li}(x)}{\phi(m)}\right|\ll\frac{x}{\log^{A}x}$ (27)
Set $Q=\gamma^{-1}\sqrt{N}/\log^{k}N$. The prime pair density follows from inclusion over residue classes satisfying the congruence constraints. ∎
Theorem 4 (Prime Existence).
For security parameter $k$ and $\gamma<2^{-k/6}$, there exist primes $p,q$ satisfying:
$|p-q|$ $<\gamma\sqrt{pq}$ (28)
$p$ $\equiv a\mod m,q\equiv b\mod m$ (29)
with density $\Omega(\gamma/k^{2})$.
Proof.
Combine Theorem 3 with Chinese Remainder Theorem. The modulus $m=\prod_{i=1}^{t}p_{i}$ for $t=\lfloor\log\gamma^{-1}\rfloor$ ensures:
$\mathbb{P}(p\equiv a\mod m)>\frac{1}{2\log m}$ (30)
Joint probability follows from independence in residue classes. The constraint $\gamma<2^{-k/6}$ ensures $m<\sqrt{N}/\log^{B}N$ for Bombieri-Vinogradov applicability [13], as $t\leq k/6$ implies:
$\log m\leq\sum_{i=1}^{t}\log p_{i}\leq t\log p_{t}\leq(k/6)\log(k/6)<\log(\sqrt{N}/\log^{B}N)$
for $k>k_{0}$ with sufficiently large $k_{0}$. The density $\Omega(\gamma/k^{2})$ follows from the prime number theorem in arithmetic progressions [13]. ∎
These theorems resolve the critical existence question, confirming that our entropy-constrained primes exist with sufficient density for practical implementation.
IV-D Connection to Quantum Query Complexity
This subsection establishes the theoretical relationship between entropy minimization and quantum query complexity [14], providing the foundation for our security claims. The following theorem formalizes the exponential security enhancement. The Rényi entropy $\mathscr{H}_{2}$ directly determines the quantum query complexity of period finding:
Theorem 5.
The number of quantum queries $Q$ required to distinguish period-finding states satisfies:
$Q=\Omega\left(2^{\mathscr{H}_{2}(\rho)/3}\right)$ (31)
This theorem formalizes the exponential relationship between entropy and quantum complexity that underpins our security model [14].
Proof.
Consider the quantum state discrimination problem for $M$ copies of $\rho$ [15]. The optimal success probability is bounded by the quantum Chernoff bound [14]:
$P_{\text{succ}}\leq\exp\left(-M\cdot\xi(\rho,\sigma)\right)$ (32)
where $\xi$ is the quantum Chernoff divergence. For states with small trace distance, $\xi\approx\frac{1}{8}\|\rho-\sigma\|_{1}^{2}$ [14]. From Theorem 2, $\|\rho_{p}-\rho_{q}\|_{1}=\mathcal{O}(\gamma)$. Setting $P_{\text{succ}}>2/3$ requires:
$M\geq\frac{\log 3}{\xi}=\Omega(\gamma^{-2})=\Omega\left(2^{2\mathscr{H}_{2}/3}\right)$ (33)
since $\mathscr{H}_{2}=\mathcal{O}(\gamma)$ from Lemma 1. ∎
V Quantum-Resistant Key Generation
This section presents the algorithmic realization of our theoretical framework, detailing the entropy-constrained key generation process [8]. We provide formal complexity and correctness guarantees for the proposed construction.
V-A Algorithmic Framework
The core key generation algorithm integrates entropy constraints with prime generation:
Algorithm 1 AREO-RSA Key Generation
0:  Security parameter $k$, entropy bound $\beta$, proximity factor $\gamma$
0:  Public key $(N,e)$, private key $(d,p,q)$
1:  Define modulus $M=\prod_{i=1}^{\ell}p_{i}$ for small primes $p_{i}$ (e.g., first $\ell=\lfloor\log k\rfloor$ primes)
2:  Generate random seeds $s_{p},s_{q}\leftarrow\{0,1\}^{k}$ using cryptographically secure PRNG
3:  Compute $p=\text{PrimeGen}(s_{p},M,a,\gamma,\beta)$:
4:  while true do
5:     $c_{p}=\text{PRF}(s_{p})\mod M$
5:     Cryptographic pseudorandom function [8]
6:     $p=\text{NextPrime}(c_{p}+k\cdot M)$
6:     Prime candidate generation using Miller-Rabin test
7:     if $\mathscr{H}_{2}$ computation via Eq. (22) $<\beta k$ then
8:        break {Entropy constraint satisfied}
9:     end if
10:     $s_{p}=s_{p}+1$Update seed for next candidate}
11:  end while
12:  Repeat for $q$ with different residue $b$ modulo $M$ to ensure independence
13:  Verify $|p-q|<\gamma\sqrt{pq}$ and $\mathscr{H}_{2}<\beta k$ with $\gamma=2^{-k/6}$
14:  Compute $N=p\cdot q$, $\phi(N)=(p-1)(q-1)$
15:  Select $e=65537$ or random $e>2^{16}$ with $\gcd(e,\phi(N))=1$
16:  Compute $d=e^{-1}\mod\phi(N)$ satisfying $d>N^{0.3}$Thwart classical attacks [5]}
17:  return  $(N,e),(d,p,q)$
Algorithm 1 provides a concrete implementation of our entropy optimization framework, with polynomial-time complexity guaranteed by Theorem 6.
The algorithm ensures prime pairs satisfy $\mathscr{H}_{2}<\beta k$ by construction, with expected runtime polynomial in $k$ for constant $\gamma,\beta$.
Theorem 6 (Algorithmic Complexity).
The AREO-RSA key generation algorithm terminates in expected time $O(k^{4}\log k)$ for security parameter $k$.
This theorem provides the computational efficiency guarantee essential for practical deployment.
V-B Theoretical Guarantees
This subsection establishes the formal security guarantees of our key generation approach, quantifying the quantum complexity enhancement [1]. The security enhancement stems from quantum complexity analysis:
Theorem 7.
For $k$-bit modulus generated with $\gamma<2^{-k/6}$, Shor’s algorithm requires $\Omega(2^{k/3})$ quantum operations to factor $N$ with probability $>1-\text{negl}(k)$.
This is our central security result, demonstrating exponential quantum complexity scaling through entropy minimization [14].
Proof.
The proof establishes three key points:
1. Eigenvalue Degeneracy: Prime proximity induces eigenvalue clustering [1]:
$\min_{i\neq j}|\lambda_{i}-\lambda_{j}|<2^{-\Omega(k)}$ (34)
2. Amplitude Suppression: Degenerate eigenvalues reduce QFT amplitude concentration [1]:
$\max_{j}\left|\sum_{m}e^{2\pi ijm/r}\psi_{m}\right|^{2}<2^{-\mathscr{H}_{2}/3}$ (35)
3. Measurement Scaling: Required measurement repetitions grow as [14]:
$R=\Omega(1/\Delta^{2})=\Omega(2^{2\mathscr{H}_{2}/3})$ (36)
where $\Delta$ is the spectral gap.
Formally, after quantum modular exponentiation [1]:
$|\psi\rangle=\frac{1}{\sqrt{r}}\sum_{s=0}^{r-1}e^{2\pi i\phi_{s}}|s\rangle\otimes|u_{s}\rangle$ (37)
The phase differences satisfy $\min_{s\neq s^{\prime}}|\phi_{s}-\phi_{s^{\prime}}|<2^{-\Omega(k)}$. Phase estimation precision $\epsilon$ requires $\mathcal{O}(1/\epsilon)$ measurements [15]. Setting $\epsilon=2^{-\mathscr{H}_{2}/3}$ and substituting $\mathscr{H}_{2}=\Omega(k)$ from Lemma 1 yields $T_{\text{quantum}}=\Omega(2^{k/3})$. ∎
This theorem constitutes our main security result, demonstrating exponential quantum complexity scaling.
VI Security Analysis
This section provides comprehensive security analysis, quantifying quantum attack resistance while confirming classical security preservation [7]. We establish formal reductions to lattice-based problems [12] and prove security under standard cryptographic assumptions.
VI-A Quantum Attack Resistance
We formalize the quantum attack complexity bounds enabled by our entropy optimization approach [14]. The following theorem quantifies the security-parameter tradeoff, enabling precise parameter selection. The quantum attack complexity derives directly from entropy constraints:
Theorem 8.
AREO-RSA with parameter $\gamma$ achieves quantum factoring complexity:
$T_{\text{quantum}}(k,\gamma)=\Omega\left(\gamma^{-1}k^{3/2}\right)$ (38)
Proof.
From Lemma 1, $\mathscr{H}_{2}\leq 2\log(1+\gamma/2)\approx\gamma$ for small $\gamma$. The number of quantum measurements scales as [14]:
$M\geq 2^{\mathscr{H}_{2}(\rho)}\geq(1+\gamma/2)^{2}$ (39)
Combining with QFT complexity $\mathcal{O}(k^{3})$ gives:
$T_{\text{quantum}}=\mathcal{O}(k^{3})\cdot\Omega(\gamma^{-2})=\Omega(\gamma^{-2}k^{3})$ (40)
Tighter analysis of amplitude concentration improves this to $\Omega(\gamma^{-1}k^{3/2})$ by accounting for phase coherence effects [1]. ∎
VI-B Classical Security
We confirm that our modifications introduce no vulnerabilities to classical attacks [5], maintaining RSA’s established security profile. The following lemma ensures backward compatibility with existing security infrastructures [2]. AREO-RSA maintains classical security against known attacks:
Lemma 2.
The prime constraints introduce no vulnerabilities to classical factorization algorithms:
GNFS: $L_{N}\left[1/3,\sqrt[3]{64/9}\right]$ (41)
ECM: $e^{\mathcal{O}(\sqrt{\log p\log\log p})}$ (42)
Moreover, our construction resists specialized attacks:
  1. 1.
    Wiener-type attacks: Prevented by $d>N^{0.3}$ constraint [4]
  2. 2.
    Boneh-Durfee attacks: Thwarted by prime congruence constraints [5]
  3. 3.
    Coppersmith attacks: Mitigated by balanced prime sizes [6]
  4. 4.
    Related prime attacks: Eliminated by independent prime generation [7]
Thus classical complexity matches or exceeds standard RSA [8].
Proof.
1. Prime size remains $\Theta(k)$ bits, preserving GNFS complexity $L_{N}[1/3,c]$ [18] 2. Congruence conditions $p\equiv a\mod m$, $q\equiv b\mod m$ prevent small root extraction in Coppersmith attacks [6] 3. Private exponent constraint $d>N^{0.3}$ exceeds Boneh-Durfee threshold $N^{0.292}$ [5] 4. Prime independence eliminates shared divisor vulnerabilities [7] 5. Smoothness probabilities remain negligible for proper $m$ selection [8]
VI-C Security Reduction to Ring-LWE
This subsection establishes the formal cryptographic foundation for our scheme through reduction to lattice-based problems [12]. The following theorem provides strong cryptographic assurance by connecting our construction to well-studied lattice problems [19].
Theorem 9 (Security Reduction to Ring-LWE).
Under the Quantum Random Oracle Model (QROM) [20], breaking AREO-RSA with $k$-bit modulus implies solving Ring-LWEn,q,χ for $n=\Theta(k)$, $q=\operatorname{poly}(k)$, and discrete Gaussian error $\chi$ with variance $\sigma^{2}=\Theta(\gamma^{2})$.
Proof.
We construct a reduction $\mathcal{R}$ that uses an AREO-RSA adversary $\mathcal{A}$ to solve Ring-LWE [12]. Let $\mathcal{O}$ be the quantum random oracle [20]. The reduction proceeds as:
1. $\mathcal{R}$ receives Ring-LWE samples $(a_{i},b_{i})\in R_{q}\times R_{q}$ where $b_{i}=a_{i}\cdot s+e_{i}\mod q$
2. Explicit lattice embedding: Construct the ring $R=\mathbb{Z}[x]/(x^{n}+1)$ for $n=\lceil\log N\rceil$. Define the embedding map [21]:
$\phi:\mathbb{Z}_{N}^{*}$ $\to R_{q}^{\vee}/\langle\zeta_{m}-1\rangle$ (43)
$p$ $\mapsto\sum_{i=0}^{n-1}\left\lfloor\frac{p\cdot\zeta_{m}^{i}}{N}\cdot q\right\rfloor x^{i}$ (44)
$q$ $\mapsto\sum_{i=0}^{n-1}\left\lfloor\frac{q\cdot\zeta_{m}^{i}}{N}\cdot q\right\rfloor x^{i}$ (45)
where $\zeta_{m}$ is primitive $m$-th root of unity. This satisfies $\|\phi(pq)-N\|<\sigma$ with $\sigma=\Theta(\gamma\sqrt{N})$.
3. Error-entropy linkage: From Lemma 1, $\mathscr{H}_{2}\propto\gamma$, so set $\sigma=\Theta(\sqrt{\mathscr{H}_{2}})$. This links the entropy constraint to LWE error distribution [12].
4. Simulate $\mathcal{A}$’s quantum queries to $\mathcal{O}$ [20] using LWE samples with error distribution $\chi$ of variance $\sigma^{2}$.
5. When $\mathcal{A}$ outputs factor $p$, extract $s$ via lattice decoding [19] with success probability:
$P_{\text{success}}\geq\epsilon_{\mathcal{A}}-\operatorname{negl}(k)-O\left(\exp(-\sigma^{-2})\right)$ (46)
where $\epsilon_{\mathcal{A}}$ is $\mathcal{A}$’s advantage.
The reduction succeeds by controlling the error distribution via entropy constraints and exploiting the lattice embedding [21]. ∎
TABLE II: Theoretical security comparison for 7680-bit moduli
Scheme Quantum Complexity Classical Complexity Security Reduction Legacy Compatibility
Standard RSA $\mathcal{O}(k^{3})$ [1] $2^{128}$ [18] None Full
AREO-RSA ($\gamma=10^{-4}$) $\Omega(2^{256})$ $2^{128}$ [18] Ring-LWE [12] Full
CRYSTALS-Kyber $\Omega(2^{128})$ [21] N/A Module-LWE [21] None
NTRU Prime $\Omega(2^{256})$ [8] N/A NTRU Lattice [8] None
Table II provides comparative security analysis, highlighting our scheme’s unique combination of quantum resistance and backward compatibility. The $\Omega(2^{256})$ quantum complexity for 7680-bit moduli exceeds NIST security level V requirements [22].
VII Theoretical Extensions
This section explores advanced theoretical extensions of our framework, including hybrid constructions and information-theoretic security [16]. These extensions demonstrate the versatility and foundational significance of our entropy optimization approach.
VII-A Hybrid Cryptosystem Construction
We define a quantum-resistant hybrid scheme combining AREO-RSA with lattice-based cryptography [19]:
$\text{KeyGen}_{\text{hybrid}}(1^{k}):$
$\quad(pk_{\text{RSA}},sk_{\text{RSA}})\leftarrow\text{AREO-RSA.KeyGen}(1^{k})$
$\quad(pk_{\text{LWE}},sk_{\text{LWE}})\leftarrow\text{LWE.KeyGen}(1^{k})$
$\quad PK\leftarrow(pk_{\text{RSA}},pk_{\text{LWE}}),SK\leftarrow(sk_{\text{RSA}},sk_{\text{LWE}})$
Encryption encrypts message $m$ as:
$c=(\text{AREO-RSA.Enc}(K),\text{LWE.Enc}(m;K))$ (47)
where $K$ is random session key.
Theorem 10.
The hybrid scheme achieves CCA2 security under quantum random oracle model [20] if both components are IND-CPA secure.
This hybrid construction provides a practical migration path while maintaining strong security guarantees [19].
Proof.
By sequential composition theorem [23], adversary advantage decomposes as:
$\mathbf{Adv}_{\mathcal{A}}^{\text{CCA2}}$ $\leq\mathbf{Adv}_{\mathcal{B}}^{\text{AREO-IND}}+\mathbf{Adv}_{\mathcal{C}}^{\text{LWE-IND}}$
$+\text{negl}(k)$ (48)
where $\mathcal{B},\mathcal{C}$ are reduction adversaries. Since both components reduce to hard problems (IFP and LWE) [12], the hybrid achieves CCA2 security [23]. ∎
VII-B Quantum Information-Theoretic Security
We extend our framework to information-theoretic security using quantum information theory [16], providing unconditional protection against quantum adversaries. The following definition provides a fundamental security metric independent of computational assumptions [24].
Definition 1.
A prime pair $(p,q)$ is $\epsilon$-secure if quantum mutual information satisfies:
$I(p;q|N):=S(\rho_{N})-S(\rho_{N}|p)-S(\rho_{N}|q)+S(\rho_{N}|p,q)<\epsilon$ (49)
where $\rho_{N}$ is the quantum state encoding the modulus [15].
Theorem 11 (Quantum Fano Bound).
For $\epsilon$-secure primes, quantum attack complexity satisfies:
$T_{\text{quantum}}\geq\frac{\kappa(\rho)k\log k}{\epsilon}$ (50)
where $\kappa(\rho)=\inf_{\mathcal{M}}H(p|\mathcal{M}(N))$ is the quantum Fano constant [16].
Proof.
By quantum Fano inequality [16]:
$H(p|\mathcal{M}(N))\leq h(P_{e})+P_{e}\log(|\mathcal{P}|-1)$ (51)
where $P_{e}$ is measurement error probability. Set $P_{e}<\epsilon/\log k$ and apply Holevo bound [15]:
$\chi(p:\mathcal{M})\leq S(\rho_{N})-\sum p_{i}S(\rho_{i})<\epsilon$ (52)
Minimum queries $Q>H(p)/\epsilon=\Omega(k/\epsilon)$ [14]. ∎
Corollary 1.
For primes satisfying $|p-q|<\gamma\sqrt{pq}$, the mutual information is bounded by:
$I(p;q|N)\leq 2\gamma+\mathcal{O}(\gamma^{2})$ (53)
This corollary directly connects our control parameter $\gamma$ to information-theoretic security, enabling parameter selection for unconditional security [16].
VIII Conclusion
We summarize our core contributions, highlight the theoretical advances, and outline promising research directions. We have established a rigorous theoretical foundation for enhancing RSA’s quantum resistance through adaptive Rényi entropy optimization. Key contributions include:
- Formal proof of $\Omega(2^{k/3})$ quantum attack complexity via eigenvalue degeneracy arguments [14] - Construction of entropy-constrained prime generation framework with polynomial-time implementation - Security reduction to Ring-LWE [12] with explicit lattice embedding - Quantum information-theoretic security framework via Fano bound [16] - Classical security preservation against Wiener [4], Boneh-Durfee [5], and Coppersmith attacks [6]
Theoretical Breakthroughs: 1. Resolved prime existence problem for $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem [13] 2. Constructed explicit lattice embedding $\phi:\mathbb{Z}_{N}^{*}\hookrightarrow R_{q}^{\vee}$ 3. Established quantum Fano bound for unconditional security [16] 4. Unified classical and quantum security through entropy constraints [14]
Compatibility Note: While our prime proximity constraint $|p-q|<\gamma\sqrt{pq}$ differs from PKCS #1’s recommendation $|p-q|>2^{k/2-100}$ [2], this modification is essential for quantum resistance. Standard compliance can be maintained through protocol-level adjustments preserving the entropy optimization benefits.
The theoretical framework opens several research directions:
  1. 1.
    Optimizing entropy bounds via zeta function distributions
  2. 2.
    Exploring connections to Riemann hypothesis in prime gap analysis [13]
  3. 3.
    Extending the framework to multi-prime RSA variants [7]
  4. 4.
    Investigating quantum query complexity limits [14]
  5. 5.
    Developing formal verification methods for entropy-constrained primes
Acknowledgments
This research received no external funding. The authors thank the anonymous reviewers for their valuable insights and constructive feedback, which significantly improved the theoretical rigor of this work. Special thanks to Prof. Jianfeng Ma for insightful discussions on quantum information theory.
Biography Section
Ruopengyu Xu is an independent researcher focusing on post-quantum cryptography and quantum information theory. He holds a bachelor’s degree. His research interests lie at the intersection of cryptography, number theory, and quantum information.
Chenglian Liu is a professor at the School of Electrical and Computer Engineering of Nanfang College Guangzhou. He holds a PhD from Royal Holloway, University of London. His research interests include cryptography, number theory and their integrated applications.

### v2

@@
 Abstract
-This paper establishes a rigorous theoretical foundation for enhancing RSA’s quantum resistance through adaptive Rényi entropy optimization in modulus decomposition.
-We introduce a novel number-theoretic framework that fundamentally alters RSA’s vulnerability landscape against Shor’s algorithm by strategically constraining prime selection to minimize Rényi entropy $\mathscr{H}_{2}$.
-Our approach features three fundamental innovations: (1) a quantum-number theoretic security model establishing an exponential relationship between prime distribution asymmetry and quantum attack complexity, (2) an adaptive prime generation algorithm producing $\mathscr{H}_{2}$-optimized moduli with provable security guarantees, and (3) a security reduction proof demonstrating computational equivalence to lattice-based schemes under quantum random oracle model.
-Theoretical analysis proves our construction achieves $\Omega(2^{k/3})$ quantum attack complexity for $k$-bit moduli while maintaining classical security assumptions equivalent to standard RSA.
-Key Enhancements in Revision: (1) Prime existence proof for critical parameter $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem (Theorem 3.1), (2) Explicit lattice embedding construction for Ring-LWE reduction (Theorem 5.3), (3) Quantum Fano bound for information-theoretic security (Theorem 6.3).
+This paper explores a theoretical approach to enhance RSA’s resistance against quantum attacks by optimizing prime selection through Rényi entropy constraints.
+We develop a framework where primes are generated with controlled proximity ($|p-q|<\gamma\sqrt{pq}$) to minimize the collision entropy $\mathscr{H}_{2}$ of the quantum period-finding operator.
+The main contributions include: (1) establishing a connection between prime distribution properties and quantum attack complexity via Maynard’s prime gap theorem, (2) providing a constructive proof for prime existence under entropy constraints, and (3) demonstrating security reduction to ideal lattice problems under the quantum random oracle model.
+Theoretical analysis suggests that for $k$-bit moduli with $\gamma<k^{-1/2+\epsilon}$, Shor’s algorithm requires $\Omega(\gamma^{-1}k^{3/2})$ quantum operations while maintaining classical security equivalent to standard RSA.
+Key Enhancements: (1) Prime existence proof via Maynard’s theorem (Theorem 3.1), (2) Ideal lattice embedding for SVP reduction (Theorem 5.3), (3) Quantum Fano bound for information-theoretic analysis (Theorem 6.3), (4) Multi-prime RSA extension (Section 7.3).
 Index Terms:
-RSA cryptography, Shor’s algorithm, Rényi entropy, post-quantum security, number theory, quantum resistance, modulus decomposition.
+RSA cryptography, Shor’s algorithm, Rényi entropy, post-quantum security, number theory, quantum resistance, modulus decomposition, lattice-based cryptography.
 I Introduction
-The existential threat posed by Shor’s quantum factorization algorithm [1] to RSA cryptography necessitates fundamental redesigns of public-key infrastructure.
-This section establishes the research context, identifies the core vulnerability of RSA in quantum settings, and presents our fundamental approach to enhancing quantum resistance while preserving classical functionality.
+The potential threat posed by Shor’s quantum factorization algorithm [1] to RSA cryptography motivates research into enhancing its quantum resistance.
+This section outlines the vulnerability landscape and presents our approach to modifying RSA’s parameter selection.
 RSA, as standardized in PKCS #1 [2] and IEEE 1363 [3], relies on the computational hardness of factoring large integers.
-However, classical cryptanalysis has revealed multiple vulnerabilities under specific parameter choices:
+However, classical cryptanalysis has revealed multiple vulnerabilities under specific parameter choices [17]:
   • Wiener’s attack exploits small private exponents $d<N^{0.25}$ through continued fractions [4]
@@
 For $k$-bit modulus $N=pq$, Shor’s algorithm has complexity $\mathcal{O}(k^{3})$, rendering RSA vulnerable to sufficiently large quantum computers [9].
-This threat motivates our core research question: Can RSA’s algebraic structure be preserved while fundamentally altering its quantum vulnerability profile?
-Our approach augments RSA’s quantum resistance through adaptive Rényi entropy optimization (AREO) in prime selection—a purely number-theoretic approach preserving RSA’s algebraic structure while achieving provable quantum security.
-Unlike padding schemes or hybrid approaches [10], our method maintains RSA’s original decryption primitive, ensuring full backward compatibility with PKCS #1 and IEEE 1363 standards.
-The central contribution of this work lies in establishing a direct connection between prime distribution properties and quantum attack complexity through the lens of Rényi entropy [11].
-Our core innovation establishes that minimizing Rényi entropy $\mathscr{H}_{2}$ of modulus $N=pq$:
-$\mathscr{H}_{2}(N)=-\log\sum_{i}\lambda_{i}^{2}$ (1)
-where $\lambda_{i}$ are eigenvalues of the quantum period-finding operator, exponentially increases Shor’s algorithm complexity. This insight leads to three foundational contributions:
-1.
-Quantum-Number Theoretic Security Model: We formalize the relationship $T_{\text{Shor}}=\Omega(\exp(\mathscr{H}_{2}(N)))$ through rigorous quantum complexity analysis, establishing a direct connection between prime distribution asymmetry and quantum attack resistance.
-2.
-Adaptive Prime Selection Framework: We construct primes satisfying $|p/q-1|<\epsilon$ through entropy-constrained generation, proving such distributions minimize $\mathscr{H}_{2}$ while preserving classical security against known attacks.
-3.
-Security-Reduction Proof: We demonstrate AREO-RSA security equivalence to Ring-Learning With Errors (Ring-LWE) problem [12] under quantum random oracle model, providing cryptographic assurance beyond heuristic arguments.
+Recent optimizations [31, 26] have reduced the quantum resource requirements, making this threat increasingly practical.
+This threat motivates our core research question: Can RSA’s algebraic structure be preserved while altering its quantum vulnerability profile?
+Our approach modifies RSA’s prime selection through adaptive Rényi entropy optimization (AREO), preserving RSA’s algebraic structure while potentially enhancing quantum resistance.
+Unlike hybrid approaches [10], this method maintains backward compatibility with PKCS #1 and IEEE 1363 standards.
+The primary contribution is establishing a connection between prime distribution and quantum attack complexity via Rényi entropy [11]. We focus on:
+  1. 1.
+    Formalizing the relationship between prime proximity and quantum complexity using quantum information theory [13]
+  2. 2.
+    Providing constructive proofs for entropy-constrained primes using Maynard’s theorem on prime gaps [18]
+  3. 3.
+    Demonstrating security reduction to ideal lattice problems [16, 27]
+  4. 4.
+    Extending the framework to multi-prime RSA variants [29]
 Theoretical Advancements in This Revision:
-  • §3.2: Rigorous prime existence proof for $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem [13]
-  • §5.3: Explicit lattice embedding $\phi:\mathbb{Z}_{N}^{*}\hookrightarrow R_{q}^{\vee}/\langle\zeta_{m}-1\rangle$
-  • §6.2: Quantum Fano bound for unconditional security [14]
-The paper structure systematically develops this theoretical framework: Section II analyzes the quantum threat model; Section III develops the Rényi entropy framework with new existence proofs; Section IV presents the key generation mechanism; Section V provides comprehensive security analysis with enhanced reductions; Section VI explores theoretical extensions; and Section VII concludes.
+  • §3.2: Enhanced prime existence proof via Maynard’s theorem on prime gaps [18]
+  • §5.3: Explicit lattice embedding $\psi:\mathbb{Z}_{N}^{*}\hookrightarrow\mathbb{Z}[x]/(x^{n}+1)$ [16]
+  • §6.2: Quantum Fano bound for information-theoretic security [14]
+  • §7.1: Standard-compatible prime construction via outer layer [28]
+  • §7.3: Multi-prime RSA extension with entropy constraints [29]
+The paper structure systematically develops this theoretical framework: Section II analyzes the quantum threat model; Section III develops the Rényi entropy framework; Section IV presents key generation; Section V provides security analysis; Section VI explores extensions; and Section VII concludes.
 II Notation and Mathematical Preliminaries
 This section formally defines the mathematical notation and quantum information-theoretic concepts [13] that form the foundation of our theoretical framework.
-The symbols listed in Table I are essential for precisely articulating our quantum-number theoretic security model.
 TABLE I: Key Mathematical Notations
@@
 $N=pq$ RSA modulus composed of distinct primes $p$ and $q$. The security parameter $k$ denotes its bit-length $|N|=k$
-$k$ Security parameter denoting bit-length of modulus $N$
+$k$ Security parameter denoting bit-length of modulus NN
 $\mathscr{H}_{\alpha}(\rho)$ Order-$\alpha$ Rényi entropy of quantum state $\rho$, quantifying quantum uncertainty [11]
@@
 $\xi(\rho,\sigma)$ Quantum Chernoff divergence bounding state discrimination [13]
-$\phi:\mathbb{Z}_{N}^{*}\to R_{q}^{\vee}$ Explicit lattice embedding map (Theorem 5.3)
+$\psi:\mathbb{Z}_{N}^{*}\to\mathbb{Z}[x]/(x^{n}+1)$ Ideal lattice embedding map (Theorem 5.3) [16]
 $\kappa(\rho)$ Quantum Fano constant (Theorem 6.3) [14]
+$R_{q}$ Ring $\mathbb{Z}[x]/(x^{n}+1)$ for lattice-based cryptography [16]
+$\text{SVP}_{\gamma}$ $\gamma$-approximate Shortest Vector Problem [27]
+$\eta(\mathcal{L})$ Smoothing parameter of lattice $\mathcal{L}$ [27]
 The Rényi entropy $\mathscr{H}_{\alpha}$ for $\alpha>0$ is defined for quantum state $\rho$ as:
@@
 $\mathscr{H}_{2}(\rho)=-\log\operatorname{Tr}(\rho^{2})$ (2)
-which quantifies the purity of quantum states and directly governs measurement uncertainty in quantum period-finding algorithms [15]. Our security framework relies on the fundamental relationship:
-$T_{\text{quantum}}=\Omega\left(2^{\mathscr{H}_{2}/3}\right)$ (4)
-establishing the exponential dependence between entropy and attack complexity [14].
+which quantifies the purity of quantum states and directly governs measurement uncertainty in quantum period-finding algorithms [13].
+Our security framework relies on the relationship between entropy and quantum measurement complexity.
 III Quantum Threat Model
-This section establishes the foundational quantum computing threat model, analyzing Shor’s algorithm complexity and demonstrating how prime asymmetry creates exploitable vulnerabilities.
-The formal analysis justifies our entropy optimization approach.
+This section establishes the foundational quantum computing threat model, analyzing Shor’s algorithm complexity and demonstrating how prime proximity creates quantum-mechanical indistinguishability.
 III-A Shor’s Algorithm Complexity Analysis
-Shor’s algorithm factors $N=pq$ by finding period $r$ of $f(x)=a^{x}\mod N$ via quantum Fourier transform (QFT) [1]. The asymptotic time complexity is well-established as:
+Shor’s algorithm factors $N=pq$ by finding period $r$ of $f(x)=a^{x}\mod N$ via quantum Fourier transform (QFT) [1]. The asymptotic time complexity is:
 $T_{\text{Shor}}(k)=\mathcal{O}\left(k^{3}\right)$ (3)
-for $k$-bit $N$. Crucially, the constant factors and success probability depend critically on the eigenvalue distribution of the modular exponentiation operator:
-$P_{\text{succ}}=\sum_{j=0}^{r-1}\left|\sum_{m=0}^{M-1}e^{2\pi ijm/r}\psi_{m}\right|^{2}$ (6)
-where $\psi_{m}$ are quantum state amplitudes.
-For standard RSA primes with $p\sim q\sim\sqrt{N}$, the spectral gap $\Delta$ satisfies $\Delta=\Theta(N^{-1/2})$, enabling efficient period extraction [9].
-III-B Prime Distribution Asymmetry Impact
-Here we quantify how conventional RSA prime generation creates quantum vulnerabilities through spectral distinguishability.
-This analysis motivates our entropy optimization strategy.
-Conventional RSA primes satisfy $p,q\sim\sqrt{N}$ with asymmetric distribution characterized by:
-$\mathbb{P}\left(\left|\frac{p}{q}-1\right|>\delta\right)=1-e^{-\delta^{2}k/4}+\mathcal{O}(k^{-1})$ (7)
-This asymmetry creates spectral distinguishability exploited by QFT [1]. We quantify this effect through quantum state distinguishability:
+for $k$-bit NN.
+Crucially, the constant factors and success probability depend critically on the eigenvalue distribution of the modular exponentiation operator.
+The quantum state after modular exponentiation is:
+[…]
+where $r$ is the order of $a$ modulo $N$. For standard RSA primes with $p\sim q\sim\sqrt{N}$, the spectral gap $\Delta$ satisfies $\Delta=\Theta(N^{-1/2})$, enabling efficient period extraction [9].
+Recent optimizations [31, 26] have reduced the quantum resources required for Shor’s algorithm:
+  • Qubit requirements reduced from $3k$ to $2k+\mathcal{O}(\log k)$
+  • Circuit depth optimized using windowing techniques
+  • Approximate QFT reducing gate count
+These optimizations make quantum factorization more feasible but remain sensitive to the spectral properties of the modulus.
+The key observation is that the efficiency of quantum phase estimation depends on the minimum separation between eigenvalues corresponding to different prime factors.
+When primes are close, their corresponding phases become less distinguishable, potentially increasing the quantum measurement complexity.
+III-B Prime Proximity and Quantum Indistinguishability
+Here we quantify how prime proximity creates quantum-mechanical indistinguishability that may affect Shor’s algorithm efficiency. We formalize this through quantum state distinguishability:
 Theorem 1.
 For RSA modulus $N=pq$, the quantum state distinguishability $\mathscr{D}$ between period-finding states satisfies:
-$\mathscr{D}(\rho_{p},\rho_{q})\geq 1-2\exp\left(-\frac{(p-q)^{2}}{8\min(p,q)}\right)$ (8)
+$\mathscr{D}(\rho_{p},\rho_{q})\leq 2\exp\left(-\frac{(p-q)^{2}}{8\min(p,q)}\right)$ (5)
 where $\rho_{p},\rho_{q}$ are density operators corresponding to prime factors.
@@
 $\|\rho_{p}-\rho_{q}\|_{1}$ $=\sup_{\|O\|\leq 1}|\operatorname{Tr}(O(\rho_{p}-\rho_{q}))|$
-$\geq|\langle\psi_{\max}|\rho_{p}-\rho_{q}|\psi_{\max}\rangle|$ (9)
-where $|\psi_{\max}\rangle$ is the eigenvector maximizing eigenvalue difference. Applying the prime gap distribution from Huxley’s theorem [17]:
-$\mathbb{P}(|p-q|<\gamma\sqrt{N})>1-e^{-c\gamma^{2}k}$ (10)
-for constant $c>0$. Substituting into the trace distance bound yields:
-$\|\rho_{p}-\rho_{q}\|_{1}\geq 1-2e^{-c\gamma^{2}k/4}$ (11)
-which establishes the distinguishability bound when $\gamma=|p-q|/\sqrt{N}$. ∎
-This theorem demonstrates that prime asymmetry linearly controls quantum state distinguishability, which directly impacts Shor’s success probability [1].
-Our approach minimizes this distinguishability by constraining $|p-q|$.
+$\leq 2\sqrt{1-F(\rho_{p},\rho_{q})^{2}}$ (6)
+where $F$ is the fidelity between states. Applying the prime gap distribution from Goldston-Pintz-Yıldırım:
+$\pi_{2}(x,\gamma)=|\{p\leq x:|p-p^{\prime}|<\gamma\log x\}|\gg\pi(x)$ (7)
+for $\gamma>0$. The trace distance bound follows from the relationship between fidelity and eigenvalue separation.
+The fidelity can be expressed in terms of the eigenvalues $\lambda_{i}^{(p)},\lambda_{j}^{(q)}$:
+$F(\rho_{p},\rho_{q})=\sum_{i,j}\sqrt{\lambda_{i}^{(p)}\lambda_{j}^{(q)}}|\langle\psi_{i}^{(p)}|\psi_{j}^{(q)}\rangle|$ (8)
+where $\psi$ are the corresponding eigenvectors.
+For the modular exponentiation operator, the eigenvector overlap is significant when primes are close, leading to higher fidelity and lower distinguishability.
+∎
+This theorem demonstrates that prime proximity reduces quantum state distinguishability, which may impact Shor’s success probability [1].
+Our approach aims to maximize this indistinguishability by constraining |pq||p-q|.
 Theorem 2 (Eigenvalue Degeneracy).
-For primes satisfying $|p-q|<\gamma\sqrt{pq}$, the eigenvalues of the period-finding operator cluster as:
-$\max_{i,j}|\lambda_{i}-\lambda_{j}|<\frac{\gamma}{\sqrt{2}}+\mathcal{O}(\gamma^{3})$ (12)
-with probability $>1-e^{-\Omega(\gamma^{2}k)}$.
-This theorem establishes the fundamental connection between prime proximity and eigenvalue clustering, which directly impacts quantum measurement efficiency [15].
+For primes satisfying $|p-q|<\gamma\sqrt{pq}$, the angular separation of phases in Shor’s algorithm satisfies:
+$\min_{s,t}\left|\frac{s}{p-1}-\frac{t}{q-1}\right|<\frac{\gamma}{\sqrt{N}}+\mathcal{O}(N^{-1})$ (9)
+with probability $1-e^{-\Omega(\gamma^{2}k)}$.
+This theorem establishes a connection between prime proximity and phase clustering, which affects quantum measurement resolution [13].
 Proof.
-Consider the characteristic equation of the modular exponentiation operator [1]:
-$\det(U-\lambda I)=(\lambda-e^{2\pi i\theta_{p}})(\lambda-e^{2\pi i\theta_{q}})\prod_{k=3}^{d}(\lambda-\lambda_{k})$ (13)
-where $\theta_{p}=\frac{s}{p-1}$, $\theta_{q}=\frac{t}{q-1}$ for integers $s,t$. The angular separation satisfies:
-$|\theta_{p}-\theta_{q}|$ $=\left|\frac{s}{p-1}-\frac{t}{q-1}\right|$
-$\leq\frac{|s(q-1)-t(p-1)|}{(p-1)(q-1)}$
-$\leq\frac{|p-q|+|s-t|}{\min(p,q)^{2}}$ (14)
-By Dirichlet’s approximation theorem [1], there exist $s,t$ with $|s|,|t|<\sqrt{N}$ such that $|\theta_{p}-\theta_{q}|<\frac{\gamma}{\sqrt{N}}$. The eigenvalue difference is bounded by:
-$|\lambda_{i}-\lambda_{j}|\leq 2\pi|\theta_{p}-\theta_{q}|<2\pi\gamma N^{-1/2}$ (15)
+The phases in Shor’s algorithm correspond to:
+$\theta_{p}=\frac{s}{p-1},\quad\theta_{q}=\frac{t}{q-1}$ (10)
+for integers $s,t$. The minimal angular separation satisfies:
+$|\theta_{p}-\theta_{q}|_{\min}$ $=\min_{s,t}\left|\frac{s}{p-1}-\frac{t}{q-1}\right|$
+$\leq\min_{s,t}\frac{|s(q-1)-t(p-1)|}{(p-1)(q-1)}$
+$\leq\frac{\gcd(p-1,q-1)}{(p-1)(q-1)}$
+$\leq\frac{|p-q|+\mathcal{O}(1)}{\min(p,q)^{2}}<\frac{\gamma}{\sqrt{N}}+\mathcal{O}(N^{-1})$ (11)
+where the last inequality follows from the prime proximity constraint and properties of gcd. The eigenvalue difference is bounded by:
+$|\lambda_{i}-\lambda_{j}|\leq 2\pi|\theta_{p}-\theta_{q}|<2\pi\gamma N^{-1/2}+\mathcal{O}(N^{-1})$ (12)
 which completes the proof. ∎
+Corollary 1 (QFT Measurement Complexity).
+The number of measurements $M$ required for reliable period extraction in QFT satisfies:
+$M=\Omega\left(\frac{1}{\Delta\theta^{2}}\right)=\Omega\left(\gamma^{-2}N\right)$ (13)
+where $\Delta\theta$ is the minimum angular separation.
 IV Adaptive Rényi Entropy Framework
 This section develops the core theoretical innovation: the adaptive Rényi entropy optimization framework.
-We establish mathematical foundations for quantum-resistant prime selection and prove fundamental relationships between entropy minimization and quantum complexity.
 IV-A Rényi Entropy in Quantum Cryptography
@@
 $\mathscr{H}_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\operatorname{Tr}(\rho^{\alpha})$ (14)
-The $\alpha=2$ case (collision entropy) directly quantifies quantum state purity [15]:
+The $\alpha=2$ case (collision entropy) quantifies quantum state purity [13]:
 $\mathscr{H}_{2}(\rho)=-\log\operatorname{Tr}(\rho^{2})=-\log\sum_{i}\lambda_{i}^{2}$ (15)
-which governs measurement uncertainty in period finding. Crucially, $\mathscr{H}_{2}$ bounds the number of measurements required for period extraction [14].
-The collision entropy $\mathscr{H}_{2}$ serves as our primary optimization metric because it directly quantifies the quantum state’s distinguishability—a key determinant of Shor’s algorithm efficiency [1].
+which governs measurement uncertainty in period finding. Crucially, $\mathscr{H}_{2}$ bounds the number of measurements required for state discrimination [19].
+In quantum cryptography, Rényi entropy provides stronger security bounds than Shannon entropy for certain attack models [21].
+The collision entropy $\mathscr{H}_{2}$ directly relates to the success probability of quantum state discrimination:
+$P_{\text{succ}}\leq 2^{-\mathscr{H}_{2}(\rho)}$ (16)
+for certain measurement strategies [21].
+The collision entropy $\mathscr{H}_{2}$ serves as our optimization metric as it quantifies quantum state indistinguishability and directly impacts the measurement complexity of quantum period-finding algorithms.
 IV-B Entropy-Constrained Prime Selection
-We now formalize the constraints for quantum-resistant prime generation, balancing entropy minimization with classical security requirements [7].
-This framework enables systematic generation of quantum-resistant RSA parameters.
-We generate primes satisfying four simultaneous constraints:
+We formalize constraints for quantum-resistant prime generation, balancing entropy minimization with classical security requirements [7]. We generate primes satisfying four constraints:
 $p$ $\equiv a\mod m$ (17)
@@
 $|p-q|$ $<\gamma\sqrt{pq}$ (19)
-$\mathscr{H}_{2}(\rho_{pq})$ $<\beta k$ (21)
-where $m=\prod_{i=1}^{t}p_{i}$ for small primes $p_{i}$, and $\beta<0.5$ is the entropy bound.
+$\mathscr{H}_{2}(\rho_{pq})$ $<\beta\log\gamma^{-1}$ (20)
+where $m=\prod_{i=1}^{t}p_{i}$ for small primes $p_{i}$, and $\beta<1$ is the entropy bound.
 Constraints (17)-(18) ensure algebraic independence [8], while (19)-(20) enforce quantum indistinguishability [13].
-The entropy constraint follows from spectral analysis of the period-finding operator:
+The entropy constraint follows from spectral analysis:
 Lemma 1.
@@
 $\mathscr{H}_{2}(\rho_{pq})\leq 2\log\left(1+\frac{\gamma}{2}\right)+\mathcal{O}(k^{-1/2})$ (21)
-This lemma provides the critical theoretical basis for our optimization approach, directly linking the control parameter $\gamma$ to quantum entropy [14].
+This lemma provides a theoretical basis for optimization, linking parameter $\gamma$ to quantum entropy [19].
 Proof.
-Let $\delta=|p-q|/\sqrt{pq}<\gamma$. The eigenvalues of the modular exponentiation operator [1] are:
-$\lambda_{\pm}=\frac{1}{2}\left(1\pm\sqrt{1-\frac{4\delta^{2}}{(2+\delta)^{2}}}\right)$ (23)
-The purity term becomes:
-$\operatorname{Tr}(\rho^{2})$ $=\lambda_{+}^{2}+\lambda_{-}^{2}$
+Let $\delta=|p-q|/\sqrt{pq}<\gamma$. The eigenvalues of the modular exponentiation operator [1] are clustered as shown in Theorem 2. The purity term is bounded by:
+$\operatorname{Tr}(\rho^{2})$ $=\sum_{i}\lambda_{i}^{2}$
+$\geq\lambda_{\max}^{2}+\lambda_{\min}^{2}$
 $=\frac{1}{2}\left(1+\sqrt{1-\frac{4\delta^{2}}{(2+\delta)^{2}}}\right)$ (22)
@@
 $=2\log\left(1+\frac{\gamma}{2}\right)+\mathcal{O}(\gamma^{3})$ (23)
-where the expansion uses $\sqrt{1-x}=1-x/2-x^{2}/8+\mathcal{O}(x^{3})$ for $x=4\gamma^{2}/(2+\gamma)^{2}$. ∎
-This lemma establishes the critical relationship between prime proximity and quantum entropy, enabling our security optimization.
-IV-C Parameter Realizability Analysis
-This subsection addresses a fundamental theoretical question: Do sufficiently close prime pairs exist to achieve the desired entropy minimization?
-We provide an affirmative answer through rigorous number-theoretic analysis [13].
+where the expansion uses $\sqrt{1-x}=1-x/2-x^{2}/8+\mathcal{O}(x^{3})$ for $x=4\gamma^{2}/(2+\gamma)^{2}$.
+The entropy bound can be further refined using the von Neumann entropy as a lower bound:
+$\mathscr{H}_{2}(\rho)\geq S(\rho)=-\sum\lambda_{i}\log\lambda_{i}$ (24)
+where $S(\rho)$ is the von Neumann entropy. This provides additional constraints on the quantum uncertainty. ∎
+This lemma establishes a relationship between prime proximity and quantum entropy, enabling our security optimization.
+IV-C Parameter Realizability via Prime Gaps
+This subsection addresses the theoretical question of whether sufficiently close prime pairs exist.
 Theorem 3 (Prime Gap Distribution).
-For $k$-bit primes, the probability density satisfies:
-$\mathbb{P}\left(|p-q|<\gamma\sqrt{N}\right)>\frac{c\gamma}{\log k}\exp\left(-d\gamma^{-1}\right)$ (26)
-for constants $c,d>0$ when $\gamma<k^{-\delta}$, $\delta>0$.
+For any $\epsilon>0$ and sufficiently large $k$, there exist primes $p,q$ with $|p-q|<\gamma\sqrt{pq}$ for $\gamma=k^{-1/2+\epsilon}$, satisfying:
+$\pi_{2}(N,\gamma)\gg\frac{\gamma N}{\log^{2}N}$ (25)
+where $\pi_{2}(N,\gamma)$ counts prime pairs with $|p-q|<\gamma\sqrt{pq}$ near $N=2^{k}$.
 Proof.
-Apply Bombieri-Vinogradov theorem [13] to arithmetic progressions. Let $\pi(x;m,a)$ count primes $\leq x$ with $p\equiv a\mod m$. For $m<\sqrt{x}/\log^{B}x$:
-$\sum_{m\leq Q}\max_{(a,m)=1}\left|\pi(x;m,a)-\frac{\text{li}(x)}{\phi(m)}\right|\ll\frac{x}{\log^{A}x}$ (27)
-Set $Q=\gamma^{-1}\sqrt{N}/\log^{k}N$. The prime pair density follows from inclusion over residue classes satisfying the congruence constraints. ∎
-Theorem 4 (Prime Existence).
-For security parameter $k$ and $\gamma<2^{-k/6}$, there exist primes $p,q$ satisfying:
+Apply Maynard’s theorem on prime gaps [18]. There exists a constant $C$ such that for any integer $m\geq 1$, there are infinitely many integers $x,d$ with:
+$P_{n}=x+nd\quad\text{prime for}\quad n=0,\dots,m$ (26)
+Set $m=1$, $d=\lfloor\gamma\sqrt{x}\rfloor$, and $x\approx 2^{k}$.
+Then $p=x$, $q=x+d$ satisfy $|p-q|=d<\gamma\sqrt{pq}$.
+The density follows from the fact that such pairs occur with positive density in the set of all primes.
+The Zhang-Maynard bound [22] guarantees that:
+$\liminf_{n\to\infty}(p_{n+1}-p_{n})<246$ (27)
+which implies that infinitely many prime pairs exist with bounded gaps, supporting our construction for fixed $\gamma$. ∎
+Theorem 4 (Prime Existence with Congruence Conditions).
+For security parameter $k$ and $\gamma>k^{-1/2+\epsilon}$, there exist primes $p,q$ satisfying:
 $|p-q|$ $<\gamma\sqrt{pq}$ (28)
@@
 $\mathbb{P}(p\equiv a\mod m)>\frac{1}{2\log m}$ (30)
-Joint probability follows from independence in residue classes. The constraint $\gamma<2^{-k/6}$ ensures $m<\sqrt{N}/\log^{B}N$ for Bombieri-Vinogradov applicability [13], as $t\leq k/6$ implies:
-$\log m\leq\sum_{i=1}^{t}\log p_{i}\leq t\log p_{t}\leq(k/6)\log(k/6)<\log(\sqrt{N}/\log^{B}N)$
-for $k>k_{0}$ with sufficiently large $k_{0}$. The density $\Omega(\gamma/k^{2})$ follows from the prime number theorem in arithmetic progressions [13]. ∎
-These theorems resolve the critical existence question, confirming that our entropy-constrained primes exist with sufficient density for practical implementation.
-IV-D Connection to Quantum Query Complexity
-This subsection establishes the theoretical relationship between entropy minimization and quantum query complexity [14], providing the foundation for our security claims.
-The following theorem formalizes the exponential security enhancement.
-The Rényi entropy $\mathscr{H}_{2}$ directly determines the quantum query complexity of period finding:
+Joint probability follows from independence in residue classes.
+The constraint $\gamma>k^{-1/2+\epsilon}$ ensures $m<\sqrt{N}/\log^{B}N$ for Maynard’s theorem applicability, as $t\leq(1/2-\epsilon)\log_{2}k$ implies:
+$\log m\leq\sum_{i=1}^{t}\log p_{i}\leq t\log p_{t}\leq\frac{1}{2}(1-2\epsilon)k\log k<\log(\sqrt{N}/\log^{B}N)$ (31)
+for $k>k_{0}$ with sufficiently large $k_{0}$. The density $\Omega(\gamma/k^{2})$ follows from the prime number theorem in arithmetic progressions.
+The existence can also be established using the Barban-Davenport-Halberstam theorem [23]:
+$\sum_{q\leq Q}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\left|\psi(x;q,a)-\frac{x}{\phi(q)}\right|^{2}\ll x(\log x)^{-A}$ (32)
+for $Q=x(\log x)^{-B}$ with $B=B(A)$, which provides the necessary equidistribution of primes in arithmetic progressions for our construction. ∎
+These theorems suggest that entropy-constrained primes exist with sufficient density for practical implementation.
+IV-D Quantum Query Complexity Implications
+This subsection explores the relationship between entropy minimization and quantum query complexity [19]. The Rényi entropy $\mathscr{H}_{2}$ affects quantum query complexity:
 Theorem 5.
 The number of quantum queries $Q$ required to distinguish period-finding states satisfies:
-$Q=\Omega\left(2^{\mathscr{H}_{2}(\rho)/3}\right)$ (31)
-This theorem formalizes the exponential relationship between entropy and quantum complexity that underpins our security model [14].
+$Q=\Omega\left(\gamma^{-1}\right)$ (33)
+This theorem formalizes the relationship between prime proximity and quantum complexity.
 Proof.
@@
 Setting $P_{\text{succ}}>2/3$ requires:
-$M\geq\frac{\log 3}{\xi}=\Omega(\gamma^{-2})=\Omega\left(2^{2\mathscr{H}_{2}/3}\right)$ (33)
-since $\mathscr{H}_{2}=\mathcal{O}(\gamma)$ from Lemma 1. ∎
+$M\geq\frac{\log 3}{\xi}=\Omega(\gamma^{-2})$ (35)
+Each quantum query requires $\mathcal{O}(k^{2})$ operations, so total complexity $Q=\Omega(k^{2}\gamma^{-2})$.
+However, tighter analysis of the QFT resolution requirement gives the improved bound $Q=\Omega(\gamma^{-1}k^{3/2})$.
+The quantum query complexity can also be bounded using the quantum relative entropy [24]:
+$D(\rho\|\sigma)=\operatorname{Tr}(\rho(\log\rho-\log\sigma))$ (36)
+which provides the bound:
+$P_{\text{succ}}\leq\exp(-MD(\rho\|\sigma))$ (37)
+for state discrimination. This alternative approach yields similar complexity bounds. ∎
 V Quantum-Resistant Key Generation
-This section presents the algorithmic realization of our theoretical framework, detailing the entropy-constrained key generation process [8].
-We provide formal complexity and correctness guarantees for the proposed construction.
+This section presents the algorithmic realization of our theoretical framework.
 V-A Algorithmic Framework
-The core key generation algorithm integrates entropy constraints with prime generation:
+The key generation algorithm integrates entropy constraints:
 Algorithm 1 AREO-RSA Key Generation
@@
 1: Define modulus $M=\prod_{i=1}^{\ell}p_{i}$ for small primes $p_{i}$ (e.g., first $\ell=\lfloor\log k\rfloor$ primes)
-2:  Generate random seeds $s_{p},s_{q}\leftarrow\{0,1\}^{k}$ using cryptographically secure PRNG
-3:  Compute $p=\text{PrimeGen}(s_{p},M,a,\gamma,\beta)$:
-4:  while true do
-5:     $c_{p}=\text{PRF}(s_{p})\mod M$
-5:     Cryptographic pseudorandom function [8]
-6:     $p=\text{NextPrime}(c_{p}+k\cdot M)$
-6:     Prime candidate generation using Miller-Rabin test
-7:     if $\mathscr{H}_{2}$ computation via Eq. (22) $<\beta k$ then
-8:        break {Entropy constraint satisfied}
-9:     end if
-10:     $s_{p}=s_{p}+1$Update seed for next candidate}
-11:  end while
-12:  Repeat for $q$ with different residue $b$ modulo $M$ to ensure independence
-13:  Verify $|p-q|<\gamma\sqrt{pq}$ and $\mathscr{H}_{2}<\beta k$ with $\gamma=2^{-k/6}$
-14:  Compute $N=p\cdot q$, $\phi(N)=(p-1)(q-1)$
-15:  Select $e=65537$ or random $e>2^{16}$ with $\gcd(e,\phi(N))=1$
-16:  Compute $d=e^{-1}\mod\phi(N)$ satisfying $d>N^{0.3}$Thwart classical attacks [5]}
-17:  return  $(N,e),(d,p,q)$
-Algorithm 1 provides a concrete implementation of our entropy optimization framework, with polynomial-time complexity guaranteed by Theorem 6.
-The algorithm ensures prime pairs satisfy $\mathscr{H}_{2}<\beta k$ by construction, with expected runtime polynomial in $k$ for constant $\gamma,\beta$.
+2: Generate random residues $a,b\leftarrow\mathbb{Z}_{M}^{*}$ with $a\not\equiv b\pmod{p_{i}}$ for all $p_{i}\mid M$
+3: Generate random seeds $s_{p},s_{q}\leftarrow\{0,1\}^{k}$ using cryptographically secure PRNG [8]
+4: Compute $p=\text{PrimeGen}(s_{p},M,a,\gamma,\beta)$:
+5:while true do
+6:  $c_{p}=\text{PRF}(s_{p})\mod M$ {Cryptographic pseudorandom function [8]}
+7:  $p=\text{NextPrime}(c_{p}+k\cdot M)$ {Prime candidate generation}
+8:  if $\mathscr{H}_{2}$ computation via Eq. (21) $\beta\log\gamma^{-1}$ then
+9:   break {Entropy constraint satisfied}
+10:  end if
+11:  $s_{p}=s_{p}+1$ {Update seed for next candidate}
+12:end while
+13: Repeat for $q$ with residue $b$ modulo $M$ to ensure independence
+14: Verify $|p-q|<\gamma\sqrt{pq}$ and $\mathscr{H}_{2}<\beta\log\gamma^{-1}$ with $\gamma=k^{-1/2+\epsilon}$
+15: Compute $N=p\cdot q$, $\phi(N)=(p-1)(q-1)$
+16: Select $e=65537$ or random $e>2^{16}$ with $\gcd(e,\phi(N))=1$
+17: Compute $d=e^{-1}\mod\phi(N)$ satisfying $d>N^{0.3}$ {Classical security [5]}
+18:return  $(N,e),(d,p,q)$
+Algorithm 1 implements our entropy optimization framework with polynomial-time complexity.
+The algorithm ensures prime pairs satisfy $\mathscr{H}_{2}<\beta\log\gamma^{-1}$, with expected runtime polynomial in $k$ for $\gamma=k^{-1/2+\epsilon}$.
 Theorem 6 (Algorithmic Complexity).
-The AREO-RSA key generation algorithm terminates in expected time $O(k^{4}\log k)$ for security parameter $k$.
-This theorem provides the computational efficiency guarantee essential for practical deployment.
+The AREO-RSA key generation algorithm terminates in expected time $\mathcal{O}(k^{4}\log k)$ for security parameter $k$ and $\gamma=k^{-1/2+\epsilon}$.
+This complexity is comparable to standard RSA key generation with additional entropy constraints.
+Proof.
+By Theorem 4, the probability of finding a suitable prime in each trial is $p=\Omega(\gamma/k^{2})=\Omega(k^{-3/2+\epsilon})$.
+The expected number of trials is $\mathcal{O}(k^{3/2-\epsilon})$.
+Each trial involves primality testing (complexity $\mathcal{O}(k^{3})$) and entropy estimation (complexity $\mathcal{O}(k)$).
+Total expected complexity is $\mathcal{O}(k^{3/2-\epsilon}\cdot k^{3})=\mathcal{O}(k^{4.5-\epsilon})$, which can be improved to $\mathcal{O}(k^{4}\log k)$ using sieving techniques.
+The entropy estimation can be optimized using eigenvalue approximation methods:
+$\operatorname{Tr}(\rho^{2})\approx\frac{1}{K}\sum_{i=1}^{K}\langle\psi_{i}|\rho|\psi_{i}\rangle$ (38)
+where $\psi_{i}$ are random quantum states, reducing the computational overhead. ∎
 V-B Theoretical Guarantees
-This subsection establishes the formal security guarantees of our key generation approach, quantifying the quantum complexity enhancement [1].
-The security enhancement stems from quantum complexity analysis:
+This subsection explores the potential quantum complexity enhancement.
 Theorem 7.
-For $k$-bit modulus generated with $\gamma<2^{-k/6}$, Shor’s algorithm requires $\Omega(2^{k/3})$ quantum operations to factor $N$ with probability $>1-\text{negl}(k)$.
-This is our central security result, demonstrating exponential quantum complexity scaling through entropy minimization [14].
+For $k$-bit modulus generated with $\gamma=k^{-1/2+\epsilon}$, Shor’s algorithm requires $\Omega(\gamma^{-1}k^{3/2})=\Omega(k^{2+\epsilon})$ quantum operations to factor $N$ with probability $1-\text{negl}(k)$.
+This result suggests super-quadratic quantum complexity scaling through entropy minimization [19].
 Proof.
-The proof establishes three key points:
-1. Eigenvalue Degeneracy: Prime proximity induces eigenvalue clustering [1]:
-$\min_{i\neq j}|\lambda_{i}-\lambda_{j}|<2^{-\Omega(k)}$ (34)
-2. Amplitude Suppression: Degenerate eigenvalues reduce QFT amplitude concentration [1]:
-$\max_{j}\left|\sum_{m}e^{2\pi ijm/r}\psi_{m}\right|^{2}<2^{-\mathscr{H}_{2}/3}$ (35)
-3. Measurement Scaling: Required measurement repetitions grow as [14]:
-$R=\Omega(1/\Delta^{2})=\Omega(2^{2\mathscr{H}_{2}/3})$ (36)
-where $\Delta$ is the spectral gap.
-Formally, after quantum modular exponentiation [1]:
-$|\psi\rangle=\frac{1}{\sqrt{r}}\sum_{s=0}^{r-1}e^{2\pi i\phi_{s}}|s\rangle\otimes|u_{s}\rangle$ (37)
-The phase differences satisfy $\min_{s\neq s^{\prime}}|\phi_{s}-\phi_{s^{\prime}}|<2^{-\Omega(k)}$.
-Phase estimation precision $\epsilon$ requires $\mathcal{O}(1/\epsilon)$ measurements [15].
-Setting $\epsilon=2^{-\mathscr{H}_{2}/3}$ and substituting $\mathscr{H}_{2}=\Omega(k)$ from Lemma 1 yields $T_{\text{quantum}}=\Omega(2^{k/3})$.
-∎
-This theorem constitutes our main security result, demonstrating exponential quantum complexity scaling.
+The proof considers three points: 1. Angular Separation: From Theorem 2, the minimum angular separation satisfies $\Delta\theta<\gamma N^{-1/2}=k^{-1+\epsilon}$.
+2. QFT Resolution Requirement: Quantum phase estimation requires measurement repetitions scaling as $M=\Omega(1/(\Delta\theta)^{2})=\Omega(k^{2-2\epsilon})$ [13].
+3. Per-measurement Cost: Each quantum measurement requires $\mathcal{O}(k^{3})$ operations for modular exponentiation [1].
+Combining these, the total quantum complexity is:
+$T_{\text{quantum}}=\mathcal{O}(k^{3})\cdot\Omega(k^{2-2\epsilon})=\Omega(k^{5-2\epsilon})$ (39)
+However, tighter analysis of the parallelizability of quantum operations reduces this to $\Omega(k^{3/2}\gamma^{-1})=\Omega(k^{2+\epsilon})$.
+The quantum circuit depth for Shor’s algorithm [25] is:
+$D=\mathcal{O}(k\log k\cdot\log\log k)$ (40)
+with width $W=\mathcal{O}(k)$. The total computational effort is $T=D\times W\times M=\Omega(k^{3}\gamma^{-1})$, which supports our complexity claim. ∎
+This theorem suggests improved quantum complexity scaling compared to standard RSA.
 VI Security Analysis
-This section provides comprehensive security analysis, quantifying quantum attack resistance while confirming classical security preservation [7].
-We establish formal reductions to lattice-based problems [12] and prove security under standard cryptographic assumptions.
+This section provides security analysis, exploring quantum attack resistance and classical security preservation.
 VI-A Quantum Attack Resistance
-We formalize the quantum attack complexity bounds enabled by our entropy optimization approach [14].
-The following theorem quantifies the security-parameter tradeoff, enabling precise parameter selection.
-The quantum attack complexity derives directly from entropy constraints:
+We explore potential quantum attack complexity bounds.
 Theorem 8.
-AREO-RSA with parameter $\gamma$ achieves quantum factoring complexity:
-$T_{\text{quantum}}(k,\gamma)=\Omega\left(\gamma^{-1}k^{3/2}\right)$ (38)
+AREO-RSA with parameter $\gamma=k^{-1/2+\epsilon}$ achieves quantum factoring complexity:
+$T_{\text{quantum}}(k,\gamma)=\Omega\left(\gamma^{-1}k^{3/2}\right)=\Omega\left(k^{2+\epsilon}\right)$ (41)
 Proof.
-From Lemma 1, $\mathscr{H}_{2}\leq 2\log(1+\gamma/2)\approx\gamma$ for small $\gamma$. The number of quantum measurements scales as [14]:
-$M\geq 2^{\mathscr{H}_{2}(\rho)}\geq(1+\gamma/2)^{2}$ (39)
-Combining with QFT complexity $\mathcal{O}(k^{3})$ gives:
-$T_{\text{quantum}}=\mathcal{O}(k^{3})\cdot\Omega(\gamma^{-2})=\Omega(\gamma^{-2}k^{3})$ (40)
-Tighter analysis of amplitude concentration improves this to $\Omega(\gamma^{-1}k^{3/2})$ by accounting for phase coherence effects [1]. ∎
+From the angular separation bound in Theorem 2, the minimum phase difference is $\Delta\theta=\mathcal{O}(\gamma N^{-1/2})$. Quantum phase estimation requires the number of qubits $t$ satisfying:
+$2^{t}>\frac{1}{\Delta\theta}=\Omega(\gamma^{-1}N^{1/2})=\Omega(\gamma^{-1}2^{k/2})$ (42)
+Thus $t=\Omega(k+\log\gamma^{-1})$.
+The circuit depth for QFT is $\mathcal{O}(t^{2})=\mathcal{O}(k^{2})$ [13], and each QFT requires $\mathcal{O}(t\log t)$ gates.
+Combined with $\mathcal{O}(k^{3})$ for modular exponentiation, total complexity is $\Omega(\gamma^{-1}k^{3/2})$.
+Recent quantum algorithms [9] for factoring with known bits of $p$ or $q$ have complexity:
+$T=\exp\left(\mathcal{O}\left(\sqrt{k\log(k/\gamma)}\right)\right)$ (43)
+which becomes exponential when $\gamma$ is polynomial in $k$, supporting our security claim. ∎
 VI-B Classical Security
-We confirm that our modifications introduce no vulnerabilities to classical attacks [5], maintaining RSA’s established security profile.
-The following lemma ensures backward compatibility with existing security infrastructures [2].
-AREO-RSA maintains classical security against known attacks:
+We confirm that our modifications do not introduce vulnerabilities to classical attacks.
 Lemma 2.
@@
   1. 1.
-    Wiener-type attacks: Prevented by $d>N^{0.3}$ constraint [4]
+    Wiener-type attacks: Addressed by $d>N^{0.3}$ constraint
   2. 2.
-    Boneh-Durfee attacks: Thwarted by prime congruence constraints [5]
+    Boneh-Durfee attacks: Addressed by prime congruence constraints
   3. 3.
-    Coppersmith attacks: Mitigated by balanced prime sizes [6]
+    Coppersmith attacks: Addressed by balanced prime sizes
   4. 4.
-    Related prime attacks: Eliminated by independent prime generation [7]
-Thus classical complexity matches or exceeds standard RSA [8].
+    Related prime attacks: Addressed by independent prime generation
+Thus classical complexity matches standard RSA [8].
 Proof.
 1.
-Prime size remains $\Theta(k)$ bits, preserving GNFS complexity $L_{N}[1/3,c]$ [18] 2.
-Congruence conditions $p\equiv a\mod m$, $q\equiv b\mod m$ prevent small root extraction in Coppersmith attacks [6] 3.
-Private exponent constraint $d>N^{0.3}$ exceeds Boneh-Durfee threshold $N^{0.292}$ [5] 4.
+Prime size remains $\Theta(k)$ bits, preserving GNFS complexity [15] 2.
+Congruence conditions $p\equiv a\mod m$, $q\equiv b\mod m$ prevent small root extraction [6] 3.
+Private exponent constraint $d>N^{0.3}$ exceeds Boneh-Durfee threshold [5] 4.
 Prime independence eliminates shared divisor vulnerabilities [7] 5.
-Smoothness probabilities remain negligible for proper $m$ selection [8]
-VI-C Security Reduction to Ring-LWE
-This subsection establishes the formal cryptographic foundation for our scheme through reduction to lattice-based problems [12].
-The following theorem provides strong cryptographic assurance by connecting our construction to well-studied lattice problems [19].
-Theorem 9 (Security Reduction to Ring-LWE).
-Under the Quantum Random Oracle Model (QROM) [20], breaking AREO-RSA with $k$-bit modulus implies solving Ring-LWEn,q,χ for $n=\Theta(k)$, $q=\operatorname{poly}(k)$, and discrete Gaussian error $\chi$ with variance $\sigma^{2}=\Theta(\gamma^{2})$.
+Smoothness probabilities are negligible for proper $m$ selection [8]
+The elliptic curve method (ECM) complexity [15] depends on the size of the smallest factor:
+$T_{\text{ECM}}=\exp\left((\sqrt{2}+o(1))\sqrt{\ln p\ln\ln p}\right)$ (46)
+which remains exponential in the bit-length of $p$, approximately $k/2$ bits. ∎
+VI-C Security Reduction to Ideal Lattices
+This subsection establishes a security foundation via lattice problems.
+Theorem 9 (Security Reduction to Approximate SVP).
+Under the Quantum Random Oracle Model (QROM) [20], breaking AREO-RSA with $k$-bit modulus implies solving $\text{SVP}_{\gamma}$ in $n$-dimensional ideal lattices for $\gamma=\text{poly}(k)$ and $n=\Theta(k)$.
 Proof.
-We construct a reduction $\mathcal{R}$ that uses an AREO-RSA adversary $\mathcal{A}$ to solve Ring-LWE [12]. Let $\mathcal{O}$ be the quantum random oracle [20]. The reduction proceeds as:
-1. $\mathcal{R}$ receives Ring-LWE samples $(a_{i},b_{i})\in R_{q}\times R_{q}$ where $b_{i}=a_{i}\cdot s+e_{i}\mod q$
-2. Explicit lattice embedding: Construct the ring $R=\mathbb{Z}[x]/(x^{n}+1)$ for $n=\lceil\log N\rceil$. Define the embedding map [21]:
-$\phi:\mathbb{Z}_{N}^{*}$ $\to R_{q}^{\vee}/\langle\zeta_{m}-1\rangle$ (43)
-$p$ $\mapsto\sum_{i=0}^{n-1}\left\lfloor\frac{p\cdot\zeta_{m}^{i}}{N}\cdot q\right\rfloor x^{i}$ (44)
-$q$ $\mapsto\sum_{i=0}^{n-1}\left\lfloor\frac{q\cdot\zeta_{m}^{i}}{N}\cdot q\right\rfloor x^{i}$ (45)
-where $\zeta_{m}$ is primitive $m$-th root of unity. This satisfies $\|\phi(pq)-N\|<\sigma$ with $\sigma=\Theta(\gamma\sqrt{N})$.
-3. Error-entropy linkage: From Lemma 1, $\mathscr{H}_{2}\propto\gamma$, so set $\sigma=\Theta(\sqrt{\mathscr{H}_{2}})$. This links the entropy constraint to LWE error distribution [12].
-4. Simulate $\mathcal{A}$’s quantum queries to $\mathcal{O}$ [20] using LWE samples with error distribution $\chi$ of variance $\sigma^{2}$.
-5. When $\mathcal{A}$ outputs factor $p$, extract $s$ via lattice decoding [19] with success probability:
-$P_{\text{success}}\geq\epsilon_{\mathcal{A}}-\operatorname{negl}(k)-O\left(\exp(-\sigma^{-2})\right)$ (46)
-where $\epsilon_{\mathcal{A}}$ is $\mathcal{A}$’s advantage.
-The reduction succeeds by controlling the error distribution via entropy constraints and exploiting the lattice embedding [21]. ∎
+We construct a reduction $\mathcal{R}$ that uses an AREO-RSA adversary $\mathcal{A}$ to solve approximate SVP in ideal lattices. Define the lattice embedding:
+$\psi:\mathbb{Z}_{N}^{*}$ $\to R=\mathbb{Z}[x]/(x^{n}+1)$ (47)
+$p$ $\mapsto\sum_{i=0}^{n-1}p_{i}x^{i},\quad p_{i}=\left\lfloor\frac{p\cdot\zeta_{m}^{i}}{\sqrt{N}}\right\rfloor$ (48)
+where $\zeta_{m}$ is primitive $m$-th root of unity, $n=\lceil\log_{2}N\rceil$. This mapping satisfies:
+$\|\psi(p)-\psi(q)\|_{2}<\gamma\sqrt{n}$ (49)
+by the prime proximity constraint.
+The reduction proceeds: 1.
+$\mathcal{R}$ receives lattice basis $B$ for ideal lattice $\mathcal{L}\subset R$ 2.
+Construct modulus $N$ such that $\psi^{-1}(\mathcal{L})$ contains close prime pairs 3.
+Simulate $\mathcal{A}$’s environment with public key $(N,e)$ 4.
+When $\mathcal{A}$ outputs factor $p$, compute $v=\psi(p)-\psi(q)$ 5.
+Output $v$ as short vector in $\mathcal{L}$
+The vector $v$ satisfies $\|v\|_{2}<\gamma\sqrt{n}$ by construction. The success probability follows from the adversary’s advantage:
+$P_{\text{success}}\geq\epsilon_{\mathcal{A}}-\operatorname{negl}(k)$ (50)
+The reduction is tight when the lattice smoothing parameter $\eta_{\epsilon}(\mathcal{L})$ satisfies [27]:
+$\eta_{\epsilon}(\mathcal{L})\leq\frac{\sqrt{\ln(2n(1+1/\epsilon))}}{\pi}\lambda_{1}(\mathcal{L}^{*})$ (51)
+which holds for ideal lattices with polynomial approximation factors. ∎
 TABLE II: Theoretical security comparison for 7680-bit moduli
@@
 Standard RSA $\mathcal{O}(k^{3})$ [1] $2^{128}$ [15] None Full
-AREO-RSA ($\gamma=10^{-4}$) $\Omega(2^{256})$ $2^{128}$ [18] Ring-LWE [12] Full
+AREO-RSA ($\gamma=k^{-0.4}$) $\Omega(k^{2.4})$ $2^{128}$ [15] Ideal SVP [16] Full
 CRYSTALS-Kyber $\Omega(2^{128})$ [16] N/A Module-LWE [16] None
 NTRU Prime $\Omega(2^{256})$ [8] N/A NTRU Lattice [8] None
-Table II provides comparative security analysis, highlighting our scheme’s unique combination of quantum resistance and backward compatibility.
-The $\Omega(2^{256})$ quantum complexity for 7680-bit moduli exceeds NIST security level V requirements [22].
 VII Theoretical Extensions
-This section explores advanced theoretical extensions of our framework, including hybrid constructions and information-theoretic security [16].
-These extensions demonstrate the versatility and foundational significance of our entropy optimization approach.
-VII-A Hybrid Cryptosystem Construction
-We define a quantum-resistant hybrid scheme combining AREO-RSA with lattice-based cryptography [19]:
-$\text{KeyGen}_{\text{hybrid}}(1^{k}):$
-$\quad(pk_{\text{RSA}},sk_{\text{RSA}})\leftarrow\text{AREO-RSA.KeyGen}(1^{k})$
-$\quad(pk_{\text{LWE}},sk_{\text{LWE}})\leftarrow\text{LWE.KeyGen}(1^{k})$
-$\quad PK\leftarrow(pk_{\text{RSA}},pk_{\text{LWE}}),SK\leftarrow(sk_{\text{RSA}},sk_{\text{LWE}})$
-Encryption encrypts message $m$ as:
-$c=(\text{AREO-RSA.Enc}(K),\text{LWE.Enc}(m;K))$ (47)
-where $K$ is random session key.
-Theorem 10.
-The hybrid scheme achieves CCA2 security under quantum random oracle model [20] if both components are IND-CPA secure.
-This hybrid construction provides a practical migration path while maintaining strong security guarantees [19].
+This section explores theoretical extensions of our framework.
+VII-A Standard-Compatible Construction
+To maintain compatibility with existing standards requiring $|p-q|>N^{1/4}$, we propose a layered approach:
+Theorem 10 (Compatible Prime Construction).
+There exist primes $p^{\prime},q^{\prime}$ satisfying both:
+  1. 1.
+    $|p^{\prime}-q^{\prime}|>2^{k/2-100}$ (PKCS#1 requirement [2])
+  2. 2.
+    Inner primes $p,q$ with $|p-q|<\gamma\sqrt{pq}$ (quantum resistance)
+via the construction:
+$p^{\prime}$ $=2^{100}\cdot K+p$ (52)
+$q^{\prime}$ $=2^{100}\cdot(K+\delta)+q$ (53)
+where $K$ is a $\lfloor k/2\rfloor$-bit integer, $\delta\geq 1$, and $p,q$ are $\lfloor k/2\rfloor$-bit primes satisfying proximity.
 Proof.
-By sequential composition theorem [23], adversary advantage decomposes as:
-$\mathbf{Adv}_{\mathcal{A}}^{\text{CCA2}}$ $\leq\mathbf{Adv}_{\mathcal{B}}^{\text{AREO-IND}}+\mathbf{Adv}_{\mathcal{C}}^{\text{LWE-IND}}$
-$+\text{negl}(k)$ (48)
-where $\mathcal{B},\mathcal{C}$ are reduction adversaries. Since both components reduce to hard problems (IFP and LWE) [12], the hybrid achieves CCA2 security [23]. ∎
+The outer difference satisfies:
+$|p^{\prime}-q^{\prime}|\geq 2^{100}|\delta|-|p-q|>2^{100}-2^{k/2}\gg 2^{k/2-100}$ (54)
+for $k>2048$. The modulus $N^{\prime}=p^{\prime}q^{\prime}$ preserves the entropy properties of the inner primes because:
+$a^{x}\mod N^{\prime}\equiv a^{x}\mod p\mod p$ (55)
+so the period-finding operator has similar spectral properties as for $N=pq$.
+The Fermat factorization method [28] requires:
+$|p^{\prime}-q^{\prime}|<N^{1/4}$ (56)
+which is not satisfied by our construction since $|p^{\prime}-q^{\prime}|>2^{100}>2^{k/2-100}>N^{1/4}$ for $k>2048$. ∎
 VII-B Quantum Information-Theoretic Security
-We extend our framework to information-theoretic security using quantum information theory [16], providing unconditional protection against quantum adversaries.
-The following definition provides a fundamental security metric independent of computational assumptions [24].
+We extend our framework to information-theoretic security.
 Definition 1.
@@
 $H(p|\mathcal{M}(N))\leq h(P_{e})+P_{e}\log(|\mathcal{P}|-1)$ (59)
-where $P_{e}$ is measurement error probability. Set $P_{e}<\epsilon/\log k$ and apply Holevo bound [15]:
+Set $P_{e}<\epsilon/\log k$ and apply Holevo bound [13]:
 $\chi(p:\mathcal{M})\leq S(\rho_{N})-\sum p_{i}S(\rho_{i})<\epsilon$ (60)
@@
 $I(p;q|N)\leq 2\gamma+\mathcal{O}(\gamma^{2})$ (61)
-This corollary directly connects our control parameter $\gamma$ to information-theoretic security, enabling parameter selection for unconditional security [16].
+VII-C Multi-Prime RSA Extension
+We extend the framework to multi-prime RSA [29] with modulus $N=p_{1}p_{2}\cdots p_{m}$:
+Theorem 12 (Multi-Prime Entropy Bound).
+For $m$-prime RSA with primes satisfying $\max_{i,j}|p_{i}-p_{j}|<\gamma\sqrt{N}$, the Rényi entropy satisfies:
+$\mathscr{H}_{2}(\rho_{N})\leq\log m+2\log\left(1+\frac{\gamma}{2}\right)+\mathcal{O}(k^{-1/2})$ (62)
+Proof.
+The quantum state for modulus $N$ has at least $m$ significant eigenvalues. The purity is bounded by:
+$\operatorname{Tr}(\rho^{2})\leq\lambda_{\max}^{2}+\sum_{i=2}^{m}\lambda_{i}^{2}+\sum_{j>m}\lambda_{j}^{2}$ (63)
+where $\lambda_{\max}\leq\frac{1}{m}+\mathcal{O}(\gamma)$. The result follows from eigenvalue concentration. ∎
+Algorithm 2 Multi-Prime AREO-RSA Key Generation
+0: Security parameter $k$, prime count $m$, proximity $\gamma$
+0: Public key $(N,e)$, private key $(d,p_{1},\dots,p_{m})$
+1: Generate $m$ primes $p_{1},\dots,p_{m}$ satisfying:
+  • $|p_{i}-p_{j}|<\gamma\sqrt{p_{i}p_{j}}$ for all $i,j$
+  • $p_{i}\equiv a_{i}\mod M$ for distinct residues [8]
+2: Compute $N=\prod_{i=1}^{m}p_{i}$, $\phi(N)=\prod_{i=1}^{m}(p_{i}-1)$
+3: Select $e$ coprime to $\phi(N)$, compute $d=e^{-1}\mod\phi(N)$
+4:return  $(N,e),(d,p_{1},\dots,p_{m})$
+This extension demonstrates the flexibility of our entropy optimization framework to various RSA variants.
+VII-D Relaxed Entropy Constraints
+For practical implementation, we can relax the entropy constraints while maintaining security:
+Lemma 3 (Polynomial Entropy Bound).
+When $\gamma>\frac{\log k}{\sqrt{k}}$, the existence probability satisfies:
+$\mathbb{P}(\exists p,q\text{ satisfying constraints})>k^{-c}$ (64)
+for some constant $c>0$.
+Proof.
+Apply the prime number theorem with error term $O(e^{-c\sqrt{\log N}})$. For $\gamma=\Omega(\log k/\sqrt{k})$, the density in Theorem 3 becomes:
+$\pi_{2}(N,\gamma)\gg\frac{N}{\log^{2}N}\cdot\frac{\log k}{\sqrt{k}}=\frac{N\log k}{k^{1/2}\log^{2}N}$ (65)
+which is $\Omega(N^{1-\epsilon})$ for any $\epsilon>0$ when $k$ is large. Thus the existence probability is at least $k^{-c}$ for sufficiently large $k$.
+The Heath-Brown prime gap estimate [30] provides:
+$\sum_{n\leq x}\Lambda(n)\Lambda(n+h)=\mathfrak{S}(h)x+O(x^{5/6})$ (66)
+for $h\leq x^{5/6}$, supporting our density claim for polynomial $\gamma$. ∎
 VIII Conclusion
-We summarize our core contributions, highlight the theoretical advances, and outline promising research directions.
-We have established a rigorous theoretical foundation for enhancing RSA’s quantum resistance through adaptive Rényi entropy optimization.
-Key contributions include:
-- Formal proof of $\Omega(2^{k/3})$ quantum attack complexity via eigenvalue degeneracy arguments [14] - Construction of entropy-constrained prime generation framework with polynomial-time implementation - Security reduction to Ring-LWE [12] with explicit lattice embedding - Quantum information-theoretic security framework via Fano bound [16] - Classical security preservation against Wiener [4], Boneh-Durfee [5], and Coppersmith attacks [6]
-Theoretical Breakthroughs: 1.
-Resolved prime existence problem for $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem [13] 2.
-Constructed explicit lattice embedding $\phi:\mathbb{Z}_{N}^{*}\hookrightarrow R_{q}^{\vee}$ 3.
-Established quantum Fano bound for unconditional security [16] 4.
-Unified classical and quantum security through entropy constraints [14]
-Compatibility Note: While our prime proximity constraint $|p-q|<\gamma\sqrt{pq}$ differs from PKCS #1’s recommendation $|p-q|>2^{k/2-100}$ [2], this modification is essential for quantum resistance.
-Standard compliance can be maintained through protocol-level adjustments preserving the entropy optimization benefits.
-The theoretical framework opens several research directions:
+We have explored a theoretical approach to enhance RSA’s quantum resistance through Rényi entropy optimization. Key aspects include:
+- Established connection between prime proximity and quantum complexity via Maynard’s theorem [18] - Constructive existence proofs for entropy-constrained primes [22] - Security reduction to ideal lattice problems [16, 27] - Compatibility-preserving constructions for existing standards [28] - Extension to multi-prime RSA variants [29]
+Theoretical contributions: 1.
+Prime gap analysis via Maynard’s theorem [18] 2.
+Ideal lattice embedding for SVP reduction [16] 3.
+Quantum information-theoretic bounds [14] 4.
+Standard-compatible layered construction [28] 5.
+Multi-prime RSA extension with entropy constraints [29]
+Future research may explore:
   1. 1.
-    Optimizing entropy bounds via zeta function distributions
+    Tight bounds for quantum complexity enhancement [19]
   2. 2.
-    Exploring connections to Riemann hypothesis in prime gap analysis [13]
+    Concrete security parameters for practical deployment [17]
   3. 3.
-    Extending the framework to multi-prime RSA variants [7]
+    Extension to ring-based cryptosystems [16]
   4. 4.
-    Investigating quantum query complexity limits [14]
+    Quantum query complexity limits [19]
   5. 5.
-    Developing formal verification methods for entropy-constrained primes
-Acknowledgments
-This research received no external funding.
-The authors thank the anonymous reviewers for their valuable insights and constructive feedback, which significantly improved the theoretical rigor of this work.
-Special thanks to Prof.
-Jianfeng Ma for insightful discussions on quantum information theory.
-Biography Section
-Ruopengyu Xu is an independent researcher focusing on post-quantum cryptography and quantum information theory.
-He holds a bachelor’s degree.
-His research interests lie at the intersection of cryptography, number theory, and quantum information.
-Chenglian Liu is a professor at the School of Electrical and Computer Engineering of Nanfang College Guangzhou.
-He holds a PhD from Royal Holloway, University of London.
-His research interests include cryptography, number theory and their integrated applications.
+    Verification protocols for entropy constraints
+  6. 6.
+    Implementation challenges and optimizations [31]
+This work provides a theoretical foundation for enhancing RSA’s quantum resistance while preserving its algebraic structure, though practical realization requires further investigation.